OpenStudy (anonymous):

When you get a fraction like this: 20y/40 - 57y/40, can I swap it to 57y/40 - 20y/40?

OpenStudy (anonymous):

@kathyforver

OpenStudy (anonymous):

@kathyforever

OpenStudy (anonymous):

Yes no, maybe, I don't know

OpenStudy (anonymous):

im not completely sure. i think so but i may be wrong

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@TuringTest

OpenStudy (agent47):

One does not simply swap the fractions whenever he or she feels like it!

OpenStudy (anonymous):

AAAAAAAAAAAAARRRRRRRRRRRR

OpenStudy (agent47):

you could swap it to: -57y/50+20y/40 though

OpenStudy (anonymous):

What is 20y/40 - 57y/40?

OpenStudy (agent47):

addition is commutative, but subtraction is not.

OpenStudy (anonymous):

@headge - Do you understand negative numbers? So, in this example it's okay for the answer to be less than zero, or rewrite it as @Agent47 has done.

OpenStudy (anonymous):

Why is the 57y now over 50?

OpenStudy (markrijckenberg):

To answer your question, you need to simplify the issue. You need to put y/40 outside the brackets and the remaining numbers inside the brackets. Then you will see that 20-57 is definitely not the same value as 57-20

OpenStudy (anonymous):

Be arsed douing that!

OpenStudy (anonymous):

doing

OpenStudy (anonymous):

Mark you look like billy corgan. So I will write the question out in full

OpenStudy (anonymous):

Now *that's* an excellent example there @MarkRijckenberg @headge - I'd be careful about how I interact with the people who are helping you, ok? Now, what part of this problem don't you understand? Did Mark's example help?

OpenStudy (anonymous):

Y/2 - 5y/8 + 4y/5 = 20y/40 - 25y/40 + 32y/40 = 20y/40 - 57y/40 Then I got stuck. What do you mean? What have I said that is rude?

OpenStudy (anonymous):

Hmm, so now I'm confused. Are you supposed to solve for y in this problem? In terms of your original question, are you able to do that part? 20 - 57 = ?

OpenStudy (anonymous):

Write is as a single fraction

OpenStudy (anonymous):

And I've not been rude to Mark

OpenStudy (anonymous):

Ok, understood. So, are you able to subtract 20y from 57y? An easier way to think of it is by adding a negative, as @Agent47 did. So, -57 + 20 = ? (this answer will be a negative number FYI)

OpenStudy (anonymous):

So it could be 37y/40?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

-37y/40?

OpenStudy (anonymous):

There she goes, there she goes again. She calls my name!

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

@MarkRijckenberg

OpenStudy (anonymous):

Hehe, sorry, connection issues. But yes, you got it! It's -37y/40 :)

OpenStudy (anonymous):

I didn't know you were able to get a -y thing!

OpenStudy (anonymous):

Alright I'll give you best response

OpenStudy (markrijckenberg):

I disagree: Y/2 - 5y/8 + 4y/5 = 20y/40 - 25y/40 + 32y/40 = -5y/40 + 32y/40 = 27y/40